OpenStudy (prasadrajupvmv):

Let 'm' and 'n' be positive integers with m.n . when 8m+7n is divisible by 5, we get remainder 1 , when 7m+8n is divisible by 5 , we get remainder 2 . if m-n is divisible by 5, the remainder has to be necessatrily equal to?

OpenStudy (anonymous):

8 m + 7 n = 5 q1 + 1 7 m + 8 n = 5 q2 + 2 subtract you get m -n =5(q1-q2) -1

OpenStudy (anonymous):

What can you conclue?

OpenStudy (anonymous):

Add 5 to both sides you get m -n +5=5(q1-q2) +4 Adding 5 will not change the remainder

OpenStudy (anonymous):

When you divide m-n by 5, you get a remainder of 4

OpenStudy (prasadrajupvmv):

thnkyou itz a simple problem bt never striked into my mind

OpenStudy (anonymous):

yw